OpenStudy (anonymous):

A number, one-third of that number and one-forth of that number are added, the result is 38. what is the original nubmber?

OpenStudy (cwrw238):

n + n/3 + n/4 = 38 12n/12 + 4n/12 + 3n/12 = 38 19n/12 = 38 19n = 456 n = 456/19 = 24

OpenStudy (anonymous):

Classical algebra question. Start by calling the number \(x\). Then we can set up the equation: \[x+\frac{x}{3}+\frac{x}{4}=38\] If we multiply by 12 (to get rid of the fractions) we will get: \[12x+4x+3x=456\] Collecting the x terms: \[19x=456\] \[x=24\]